Find coefficient of correlation of the following data. Use this data to estimate y when x =


12.


X


3


6


7


10


Y


12


18


25


22



1
Expert's answer
2021-07-16T13:32:47-0400

The given data is




Regression coefficient of y on x


"b_{yx}=\dfrac{n\sum xy-\sum x.\sum y}{n\sum x^2-(\sum x)^2}"


"=\dfrac{4\times 539-26\times77}{4\times 194-(26)^2}\\[9pt]=\dfrac{154}{100}=1.54"


Regression coefficient of x on y


"b_{xy}=\dfrac{n\sum xy-\sum x.\sum y}{n\sum y^2-(\sum y)^2}"


"=\dfrac{4\times 539-26\times 77}{4\times 1577-(77)^2}\\[9pt]=\dfrac{154}{379}=0.406"



Correlation coefficient "r=\sqrt{b_{yx}\times b_{xy}}=\sqrt{1.54\times 0.406}=\sqrt{0.62}=0.787"



Here, "\bar{x}=\dfrac{\sum x}{n}=\dfrac{26}{4}=6.5"


"\bar{y}=\dfrac{\sum y}{n}=\dfrac{77}{4}=19.25"


Regression line of y on x is


"y-\bar{y}=b_{yx}(x-\bar{x})"


"\Rightarrow y-19.25=1.54(x-6.5)"


Value of y when x=12 is

"y-19.25=1.54(12-6.5)\\\Rightarrow y=8.47+19.25=27.72"